# Heavy Mesons to Charmed Tetraquark Decays

### v1

@@ VI.1 Topological amplitudes in $B^{-}_{c}\to T\overline{T}$ decays
Table 20: $B^{-}_{c}\to T\overline{T}$ decay amplitudes with $T=T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}}$ in $\Delta S=0$ transition.
# Mode $A(B^{-}_{c}\to T\overline{T})$
$(5^{\prime*},\overline{4}^{\prime*})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,-}_{\bar{c}\bar{s}}$ $\sqrt{2}C_{\bar{T}T,1}$
$(6^{\prime*},\bar{5}^{\prime*})$ $B^{-}_{c}\to T^{-}_{cs}\overline{T}^{0}_{\bar{c}\bar{s}}$ $\sqrt{2}C_{\bar{T}T,1}$
$(8^{\prime*},\bar{7}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}\bar{s}}\overline{T}^{\,--}_{\bar{c}ss}$ $2(C_{T\bar{T}}+A_{T\bar{T}})$
$(2^{\prime*},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}-2C_{T\bar{T}}-2A_{T\bar{T}})$
$(3^{\prime*},\bar{2}^{\prime*})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,-}_{\bar{c}s}$ $-\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}-2C_{T\bar{T}}-2A_{T\bar{T}})$
$(10^{\prime*},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c}\overline{T}^{\,--}_{\bar{c}}$ $\frac{2}{\sqrt{3}}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$
$(11^{\prime*},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{4}{3}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$
$(12^{\prime*},\overline{11}^{\prime*})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{0}_{\bar{c}}$ $-\frac{2}{\sqrt{3}}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$
$(3^{\prime*},\overline{15}^{\prime})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime-}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(11^{\prime*},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $-\frac{1}{3\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(12^{\prime*},\overline{14}^{\prime})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(13^{\prime},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{+}_{cd\bar{d}\bar{d}}\overline{T}^{\,--}_{\bar{c}\bar{u}dd}$ $\frac{1}{\sqrt{6}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(14^{\prime},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{3\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(15^{\prime},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(14^{\prime},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{12}(13C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(3^{\prime*},\bar{1}8^{\prime})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(12^{\prime*},\overline{17}^{\prime})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(14^{\prime},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{4\sqrt{3}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(11^{\prime*},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $-\frac{1}{\sqrt{6}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(16^{\prime},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c}\overline{T}^{\,--}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(17^{\prime},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(17^{\prime},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{4\sqrt{3}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(18^{\prime},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{2}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$
$(17^{\prime},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{4}(4C_{T}+4C_{\bar{T}}+5C_{\bar{T}T,1}+16C_{\bar{T}T,2}+2C_{T\bar{T}}+16A_{\bar{T}T}+2A_{T\bar{T}})$
@@
$(5^{\prime*},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,\prime-}_{\bar{c}}$ $-\frac{1}{2\sqrt{3}}(3C^{\prime}_{\bar{T}T,1}-2C^{\prime}_{T\bar{T}}-2A^{\prime}_{T\bar{T}})$
$(6^{\prime*},\overline{14}^{\prime})$ $B^{-}_{c}\to T^{-}_{cs}\overline{T}^{\prime 0}_{\bar{c}}$ $-\frac{1}{\sqrt{6}}(3C^{\prime}_{\bar{T}T,1}-2C^{\prime}_{T\bar{T}}-2A^{\prime}_{T\bar{T}})$
@@
$(14^{\prime},\overline{18}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $-\frac{\sqrt{3}}{4}(4C^{\prime}_{T}+C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
@@ VI.2 $B^{-}_{c}\to T\overline{T}$ decay amplitudes
The $B^{-}_{c}\to T\overline{T}$ decay amplitudes in $\Delta S=0$ and $-1$ transitions are shown in Table 19 for scenario I, and in Table 20 and 21 for scenario II. We shall concentrate on modes where both $T$ and $\overline{T}$ are flavor exotic states.
@@
Although there are three topological amplitudes $C_{T\bar{T}}$, $C_{\bar{T}T,1}$, and $A_{T\bar{T}}$ that can contribute to these modes, they occur in a unique combination, namely $C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}}$. Hence, the rates of the above two modes are highly related; in fact, they are identical
@@
This can be easily understood. $B^{-}_{c}$ is an SU(3) singlet, the operator $(\bar{c}b)(\bar{d}u)$ belongs to a ${\bf 8}$. For flavor exotic $T$ and $\overline{T}$, they form ${\bf 6}\otimes\bar{\bf 6}$, in scenario I. There is only one ${\bf 8}$ from ${\bf 6}\otimes\bar{\bf 6}={\bf 27}\oplus{\bf 8}\oplus{\bf 1}$, see Eq. (134), and, hence, there is only one combination that can match the SU(3) quantum numbers of the decaying particle $B^{-}_{c}$, the operator $(\bar{c}b)(\bar{d}u)$, and the flavor exotic final state $T$, $\overline{T}$.
@@
We now turn to scenario II. This situation is rather different. From Tables 20 and 21, we see that, we now have eight modes in $\Delta S=0$ transitions with both flavor exotic $T$ and $\overline{T}$ involved, and eight other modes in $\Delta S=-1$ transitions. Interestingly, all flavor exotic states are involved in these decays. Furthermore, the amplitudes of all these modes are linear combinations of $C^{(\prime)}_{\bar{T}T,1}$ and $C^{(\prime)}_{T\bar{T}}+A^{(\prime)}_{T\bar{T}}$. This is reasonable. Now the $T$ and $\overline{T}$ form $\overline{\bf 15}\otimes{\bf 15}$. As there are two ${\bf 8}$ from $\overline{\bf 15}\otimes{\bf 15}$, see Eq. (134) and, consequently, there are two combinations to match the SU(3) quantum numbers of the decaying particle $B^{-}_{c}$, the operator $(\bar{c}b)(\bar{d}u)$ or $(\bar{c}b)(\bar{s}u)$, and the flavor exotic final state $T$, $\overline{T}$.
@@
The final state $T$ or $\overline{T}$ can further decay to $DP$ and $DS$ or $\bar{D}P$ and $\bar{D}S$, or they can connect to these states virtually. One can obtain their full amplitudes using results in the section and those in Sec. III.
@@ VII Conclusion
In this work, we study the decays of heavy mesons to these charmed tetraquark states ($T$) using a topological amplitude approach. We first obtain the $T\to DP$ and $DS$ strong decay amplitudes by decomposing them into several topological amplitudes, where $P$ is a light psedo-scalar particle and $S$ is a low-lying scalar particle. Later on, weak decay amplitudes of $\overline{B}\to D\overline{T}$, $\overline{D}T$ and $\overline{B}\to TP$, $TS$ decays are also decomposed topologically. In addition, $B^{-}_{c}\to T\overline{T}$ decays are also discussed. Our results can be extended to other modes by replacing $D$ and $P$ with $D^{*}$ and $V$, respectively. Our main results are:
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ $\checkmark$ X 7
$\overline{B}\to D\overline{T}$, $\Delta S=0$checkmark $\checkmark$ X $\checkmark$ 7
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$checkmark $\checkmark$ $\checkmark$ X 19
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$ ()($\checkmark$)( ✓ ) ()($\checkmark$)( ✓ ) X ()($\checkmark$)( ✓ ) 19
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ $\checkmark$ X X X 8
$\overline{B}\to D\overline{T}$, $\Delta S=0$checkmark $\checkmark$ X $\checkmark$ $\checkmark$ X 8
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 20
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 21
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ X X X X 8
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 20
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$checkmark $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 21
@@
  • We introduce two different scenarios of charmed tetraquarks $T_{cq\bar{q}\bar{q}}$, where in scenario I the light antiquarks are antisymmetric, while in scenario II they are symmetric. Scenario I consists of a ${\bf 6}$ and a $\bar{\bf 3}$, while scenario II consists of a $\overline{\bf 15}$ and another $\bar{\bf 3}$. Furthermore, $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ can be in ${\bf 6}$ or $\overline{\bf 15}$, but only in one of them at the same time.
@@
  • In scenario I, we identify four flavor exotic states, namely $T^{++}_{c\bar{s}}$, $T^{+}_{c\bar{s}}$, $T^{0}_{c\bar{s}}$ and $T^{0}_{cs}$, where the first three are in the $S=+1$ isotriplet, while the last one is an $S=-1$ isosinglet, see Fig. 1 (a). They are in ${\bf 6}$. These four states are protected by their isospin and $S$ quantum numbers from mixing with other states in the scenario. Note that $T^{+}_{c\bar{s}}$ cannot be identify with $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ or $T^{*}_{c\bar{s}0}(2900)^{++}$ and it has not been considered a flavor exotic state before. It is new.
  • In scenario II, we identify twelve flavor exotic states, which are in the $S=+1$ isotriplet, $(T^{++}_{c\bar{s}},T^{+}_{c\bar{s}},T^{0}_{c\bar{s}})$, the $S=-1$ isotriplet, $(T^{+}_{cs},T^{0}_{cs},T^{-}_{cs})$, the $S=+2$ isodoublet, $(T^{++}_{c\bar{s}\bar{s}},T^{+}_{c\bar{s}\bar{s}})$, and the $S=0$ iso-quarplet, $(T_{c}^{++},T_{c}^{+},T_{c}^{0},T_{c}^{-})$, see Fig. 1 (b). They are in $\overline{\bf 15}$. These twelve states are protected by their isospin and $S$ quantum numbers from mixing with other states in the scenario. Except $T^{++}_{c\bar{s}}$, $T^{0}_{c\bar{s}}$ and $T^{0}_{cs}$, all other states cannot be identify with $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ or $T^{*}_{c\bar{s}0}(2900)^{++}$. Furthermore, $T^{+}_{c\bar{s}}$, and $T_{c}^{+}$ and $T_{c}^{0}$ have not been conseidered as flavor exotic states before.
@@
  • We decompose $T\to DP$ and $DS$ decay amplitudes into topological amplitudes, where three to four topological amplitudes are needed. However, only one topological amplitude, namely the fall-apart amplitude $F_{DP}$, is needed in $T\to DP$ decays for flavor exotic $T$ in scenarios I or II. Hence, their amplitudes are highly related. The situation in $T\to DS$ decays is similar, but with $F_{DP}$ replaced with $2C_{S}+C_{DS}$ in scenario I and $C_{DS}$ in scenario II.

### v3

@@ VI.1 Topological amplitudes in $B^{-}_{c}\to T\overline{T}$ decays
Table 20: $B^{-}_{c}\to T\overline{T}$ decay amplitudes with $T=T_{cq\{\bar{q}^{\prime}\bar{q}^{\prime\prime}\}}$ in $\Delta S=0$ transition in scenario II
. # Mode $A(B^{-}_{c}\to T\overline{T})$ $(5^{\prime*},\overline{4}^{\prime*})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,-}_{\bar{c}\bar{s}}$ $\sqrt{2}C_{\bar{T}T,1}$ $(6^{\prime*},\bar{5}^{\prime*})$ $B^{-}_{c}\to T^{-}_{cs}\overline{T}^{0}_{\bar{c}\bar{s}}$ $\sqrt{2}C_{\bar{T}T,1}$ $(8^{\prime*},\bar{7}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}\bar{s}}\overline{T}^{\,--}_{\bar{c}ss}$ $2(C_{T\bar{T}}+A_{T\bar{T}})$ $(2^{\prime*},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}-2C_{T\bar{T}}-2A_{T\bar{T}})$ $(3^{\prime*},\bar{2}^{\prime*})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,-}_{\bar{c}s}$ $\frac{1}{\sqrt{2}}(C_{\bar{T}T,1}-2C_{T\bar{T}}-2A_{T\bar{T}})$ $(10^{\prime*},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c}\overline{T}^{\,--}_{\bar{c}}$ $\frac{2}{\sqrt{3}}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$ $(11^{\prime*},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{4}{3}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$ $(12^{\prime*},\overline{11}^{\prime*})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{0}_{\bar{c}}$ $\frac{2}{\sqrt{3}}(C_{\bar{T}T,1}-C_{T\bar{T}}-A_{T\bar{T}})$ $(3^{\prime*},\overline{15}^{\prime})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime-}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(11^{\prime*},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{3\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(12^{\prime*},\overline{14}^{\prime})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(13^{\prime},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c}\overline{T}^{\,--}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(14^{\prime},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{3\sqrt{2}}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(15^{\prime},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{2}(C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(14^{\prime},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{12}(13C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(3^{\prime*},\bar{1}8^{\prime})$ $B^{-}_{c}\to T^{0}_{c\bar{s}}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{1}{2}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(12^{\prime*},\overline{17}^{\prime})$ $B^{-}_{c}\to T^{-}_{c}\overline{T}^{\prime\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(14^{\prime},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{4\sqrt{3}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(11^{\prime*},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(4C_{T}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(16^{\prime},\bar{9}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c}\overline{T}^{\,--}_{\bar{c}}$ $\frac{1}{\sqrt{2}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(17^{\prime},\overline{10}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,-}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(17^{\prime},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{4\sqrt{3}}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(18^{\prime},\bar{1}^{\prime*})$ $B^{-}_{c}\to T^{\prime\prime+}_{c\bar{s}}\overline{T}^{\,--}_{\bar{c}s}$ $\frac{1}{2}(4C_{\bar{T}}+C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}})$ $(17^{\prime},\overline{16}^{\prime})$ $B^{-}_{c}\to T^{\prime\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}}$ $\frac{1}{4}(4C_{T}+4C_{\bar{T}}+5C_{\bar{T}T,1}+16C_{\bar{T}T,2}+2C_{T\bar{T}}+16A_{\bar{T}T}+2A_{T\bar{T}})$
@@
$(5^{\prime*},\overline{13}^{\prime})$ $B^{-}_{c}\to T^{0}_{cs}\overline{T}^{\,\prime-}_{\bar{c}}$ $\frac{1}{2\sqrt{3}}(3C^{\prime}_{\bar{T}T,1}-2C^{\prime}_{T\bar{T}}-2A^{\prime}_{T\bar{T}})$
$(6^{\prime*},\overline{14}^{\prime})$ $B^{-}_{c}\to T^{-}_{cs}\overline{T}^{\prime 0}_{\bar{c}}$ $\frac{1}{\sqrt{6}}(3C^{\prime}_{\bar{T}T,1}-2C^{\prime}_{T\bar{T}}-2A^{\prime}_{T\bar{T}})$
@@
$(14^{\prime},\overline{18}^{\prime})$ $B^{-}_{c}\to T^{\prime 0}_{c}\overline{T}^{\,\prime\prime-}_{\bar{c}s}$ $\frac{\sqrt{3}}{4}(4C^{\prime}_{T}+C^{\prime}_{\bar{T}T,1}+2C^{\prime}_{T\bar{T}}+2A^{\prime}_{T\bar{T}})$
@@ VI.2 $B^{-}_{c}\to T\overline{T}$ decay amplitudes
The $B^{-}_{c}\to T\overline{T}$ decay amplitudes in $\Delta S=0$ and 1-1 transitions are shown in Table 19 for scenario I, and in Table VI.1 and 21 for scenario II. We shall concentrate on modes where both $T$ and $\overline{T}$ are flavor exotic states.
@@
Although there are three topological amplitudes $C_{T\bar{T}}$, $C_{\bar{T}T,1}$, and $A_{T\bar{T}}$ that can contribute to these modes, they occur in a unique combination, namely $C_{\bar{T}T,1}+2C_{T\bar{T}}+2A_{T\bar{T}}$. Hence, the rates of the above two modes are highly related; in fact, they are identical,
@@
This can be easily understood. $B^{-}_{c}$ is an SU(3) singlet, the operator $(\bar{c}b)(\bar{d}u)$ belongs to an eight $\bf 8$}. For flavor exotic $T$ and $\overline{T}$, they form ${\bf 6}\otimes\bar{\bf 6}$, in scenario I. There is only one $\bf 8$} from ${\bf 6}\otimes\bar{\bf 6}={\bf 27}\oplus{\bf 8}\oplus{\bf 1}$, see Eq. (176), and, hence, there is only one combination that can match the SU(3) quantum numbers of the decaying particle $B^{-}_{c}$, the operator $(\bar{c}b)(\bar{d}u)$, and the flavor exotic final state $T$, $\overline{T}$.
@@
We now turn to scenario II. This situation is rather different. From Tables VI.1 and 21, we see that, we now have eight modes in $\Delta S=0$ transitions with both flavor exotic $T$ and $\overline{T}$ involved, and eight other modes in $\Delta S=-1$ transitions. Interestingly, all flavor exotic states are involved in these decays. Furthermore, the amplitudes of all these modes are linear combinations of $C^{(\prime)}_{\bar{T}T,1}$ and $C^{(\prime)}_{T\bar{T}}+A^{(\prime)}_{T\bar{T}}$. This is reasonable. Now the $T$ and $\overline{T}$ form $\overline{\bf 15}\otimes{\bf 15}$. As there are two $\bf 8$} from $\overline{\bf 15}\otimes{\bf 15}$, see Eq. (176) and, consequently, there are two combinations to match the SU(3) quantum numbers of the decaying particle $B^{-}_{c}$, the operator $(\bar{c}b)(\bar{d}u)$ or $(\bar{c}b)(\bar{s}u)$, and the flavor exotic final state $T$, $\overline{T}$.
@@
The final state $T$ or $\overline{T}$ can further decay to $DP$ and $DS$ or $\bar{D}P$ and $\bar{D}S$, or they can connect to these states virtually. One can obtain their full amplitudes using results in this section and those in Sec. III.
@@ VII Conclusion
In this work, we study the decays of heavy mesons to these open-charmed tetraquark states ($T$) using a topological amplitude approach. We first obtain the $T\to DP$ and $DS$ strong decay amplitudes by decomposing them into several topological amplitudes, where $P$ is a light psedo-scalar particle and $S$ is a low-lying scalar particle. Subsequently, weak decay amplitudes of $\overline{B}\to D\overline{T}$, $\overline{D}T$ and $\overline{B}\to TP$, $TS$ decays are decomposed topologically. In addition, we also discuss $B^{-}_{c}\to T\overline{T}$ decays. Our results can be extended to other modes by replacing $D$ and $P$ with $D^{*}$ and $V$, respectively. Furthermore, as our analysis is based on flavor, our result can also be applied to other open-charmed tetraquark states having the same flavor structure. Our main results are:
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ $\checkmark$ X 7
$\overline{B}\to D\overline{T}$, $\Delta S=0$ $\checkmark$ $\checkmark$ X $\checkmark$ 7
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$ $\checkmark$ $\checkmark$ $\checkmark$ X 19
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$ $(\checkmark)$ $(\checkmark)$ X $(\checkmark)$ 19
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ $\checkmark$ X X X 8
$\overline{B}\to D\overline{T}$, $\Delta S=0$ $\checkmark$ $\checkmark$ X $\checkmark$ $\checkmark$ X 8
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ VI.1
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 21
@@
$\overline{B}\to D\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ X X X X 8
@@
$B^{-}_{c}\to T\overline{T}$, $\Delta S=0$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ VI.1
$B^{-}_{c}\to T\overline{T}$, $\Delta S=-1$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 21
@@
  • We introduce two different scenarios of open-charmed tetraquarks $T_{cq\bar{q}\bar{q}}$, where in scenario I the light antiquarks are antisymmetric, while in scenario II they are symmetric. Scenario I consists of a $\bf 6$} and a $\bar{\bf 3}$, while scenario II consists of a $\overline{\bf 15}$ and another $\bar{\bf 3}$. Furthermore, $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ and $T^{*}_{c\bar{s}0}(2900)^{++}$ can be in $\bf 6$} or $\overline{\bf 15}$, but can only in one of them at the same time.
@@
  • In scenario I, we identify four flavor exotic states, namely $T^{++}_{c\bar{s}}$, $T^{+}_{c\bar{s}}$, $T^{0}_{c\bar{s}}$ and $T^{0}_{cs}$, where the first three are in the $S=+1$ isotriplet, while the last one is an S=1S=-1 isosinglet, see Fig. 1 (a). They are in $\bf 6$}. These four states are protected by their isospin and $S$ quantum numbers from mixing with other states in the scenario. Note that $T^{+}_{c\bar{s}}$ cannot be identify with $T_{cs0}(2870)^{0}$, $T^{*}_{c\bar{s}0}(2900)^{0}$ or $T^{*}_{c\bar{s}0}(2900)^{++}$ and it has not been considered a flavor exotic state before. It is new.
[…]
@@
  • We decompose $T\to DP$ and $DS$ decay amplitudes into topological amplitudes, where three to four topological amplitudes are needed. However, only one topological amplitude, namely the fall-apart amplitude $F_{DP}$, is needed in $T\to DP$ decays for flavor exotic $T$ in scenarios I or II. Hence, their amplitudes are highly related. The situation in $T\to DS$ decays is similar, but with $F_{DP}$ replaced with $2C_{S}+C_{DS}$ in scenario I and $C_{DS}$ in scenario II. Although some of the modes may be kinematically forbidden, the information about their amplitudes is still useful, as the process can be a part of a larger diagram. The decomposition of decay amplitudes in terms of topological amplitudes is unaffected by rescattering. For illustration, we consider the effect of quasi-elastic rescattering on $T\to DP$ decays.